At this year's State Fair, there was a dice rolling game. If you rolled two dice and got a sum of 2 or 12, you won $10. If you rolled a 7, you won $2. Any other roll was a loss. It cost $1 to play one game with one roll of the dice. What is the expectation of the game? Round your answer to the nearest cent
Answer by greenestamps(13353)



P(2 or 12) = P(2)+P(12) = 1/36 + 1/36 = 2/36
value(2 or 12) = -1+10 = +9

P(7) = 6/36
value(7) = -1+2 = +1

P(other) = 28/36
value(other) = -1

expected value =